OpenStudy (anonymous):

What is the x-intercept of -2x + 5y = -10? A. -2 B 1/5 C 1/2 D 5 Please help me I do not know the answer

OpenStudy (anonymous):

D:5

OpenStudy (anonymous):

A: -2 \[y=\frac{2 x}{5}-2 \]

OpenStudy (anonymous):

put y=0 to find x intercept -2x+5(0)=-10 -2x+0=-10 -2x=-10 x= 10/2 x=5 s0 answer D is right answer and LagrangeSon678 also has good answer....:P

OpenStudy (anonymous):

robtobey is not doing too well tonight.

OpenStudy (anonymous):

lol